### v1

# Convergence Rates of Constrained Expected Improvement

Abstract
Constrained Bayesian optimization (CBO) methods have seen significant success in black-box optimization with constraints, and one of the most commonly used CBO methods is the constrained expected improvement (CEI) [1] algorithm. CEI is a natural extension of the expected improvement (EI)  [2] when constraints are incorporated. However, the theoretical convergence rate of CEI has not been established. In this work, we study the convergence rate of CEI by analyzing its simple regret upper bound. First, we show that when the objective function $f$ and constraint function $c$ are assumed to each lie in a reproducing kernel Hilbert space (RKHS), CEI achieves the convergence rates of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for the commonly used squared exponential and Matérn kernels, respectively. Second, we show that when $f$ and $c$ are assumed to be sampled from Gaussian processes (GPs), CEI achieves the same convergence rates with a high probability. Numerical experiments are performed to validate the theoretical analysis.
1 Introduction
Bayesian optimization (BO) is an efficient method for optimizing expensive black-box functions without derivatives. It leverages probabilistic surrogate models, most commonly Gaussian processes (GPs), to balance exploration and exploitation in the search for optimal solutions [3]. BO has found widespread success in diverse fields such as structural design [4], machine learning hyperparameter tuning [5], robotics [6], inertial confinment fusion design [7], process parameter optimization of additive manufacturing [8], etc.
While traditional BO is typically applied to unconstrained settings, many real-world problems involve black-box constraints that must be satisfied. This has motivated growing interest in constrained Bayesian optimization (CBO), where surrogate models are also constructed for constraint functions [9] that are complex and expensive to evaluate, making CBO especially valuable in applications like engineering design [10] and automated machine learning  [11]. One of the very key difference between unconstrained and constrained optimization is that the feasible region for constrained optimization problem consists of the search space where all constraints must be satisfied. A general formulation of the constrained BO problem is:
$\underset{{\bm{x}}\in C}{\text{minimize}}$ $f({\bm{x}}),$ subject to $c({\bm{x}})\leq 0,$ (1)
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is the objective function, and $c:\mathbb{R}^{d}\to\mathbb{R}^{m}$ are the constraint functions. Both are defined on a compact input space $C\subset\mathbb{R}^{d}$. The objective and the constraint functions are both expensive black-box functions, that can only be evaluated through expensive physical or computer experiments. Throughout this paper, we consider the noise-free setting for both the objective and the constraints, i.e., the function evaluations are deterministic and the true function values can be observed. In addition, a single constraint is considered, i.e., $m=1$, for simplicity of presentation. We note that our analysis can be easily extended to multiple constraints (see Remark 3.15 for details).
Broadly, CBO methods can be categorized into implicit and explicit approaches [12]. Implicit methods modify standard acquisition functions to incorporate constraints via merit functions or feasibility weights. Explicit methods estimate the feasible region directly and restrict search to this region. Among these, the constrained expected improvement (CEI) [13, 14, 1] stands out as one of the most basic and widely adopted methods. CEI is a natural extension of the well-known expected improvement (EI) function [2], where the acquisition function is computed as the product of EI and the probability of feasibility. Thanks to this simple and interpretable formulation, CEI has been successfully applied across domains, and it remains one of the default choices in many constrained BO software packages  [15].
Despite its empirical popularity, the theoretical understanding of CEI lags behind. In contrast, unconstrained EI has been more extensively studied. Under a frequentist assumption where the objective $f$ lies in a reproducing kernel Hilbert space (RKHS), [16] established the convergence rate of EI by deriving the simple regret upper bound. Other works explored the density of sampled sequences [17] or connections between EI and optimal computing budget allocation [18]. However, convergence rates (i.e., simple regret upper bound) for CEI have not been rigorously established—neither under frequentist nor under Bayesian settings. Here, Bayesian setting means the objective $f$ is a random function sampled from a GP.
Introducing constraints into EI significantly complicates the theoretical analysis. Unlike in the unconstrained case, the algorithm may need to explore infeasible regions to gain information on the constraint boundary. Furthermore, CEI’s acquisition function is inherently more complex and non-convex, posing challenges for analysis. On the other hand, the presence of constraints in CEI leads to changes in the sampling procedure. As a result, the key challenge to study the convergence rate of CEI lies in analyzing the exploration (searching for feasible regions) and exploitation (optimizing within feasible areas) since the feasibility threshold is unknown in the input space.
In this paper, we provide the first theoretical convergence rates for CEI, focusing on simple regret upper bounds under both the frequentist and Bayesian settings. Our convergence rates provide practitioners theoretical assurance for the practical deployment of CEI. Further, we would like to highlight that the obtained bounds under both settings are not direct extensions from current unconstrained results. We explain the technical challenges and how we address them in Section 3. Our contributions are summarized as follows:
  • Under the frequentist setting, we derive simple regret upper bounds of $\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)$ for the squared exponential (SE) kernel and $\mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right)$ for Matérn kernels. These bounds are improved upon the direct extension of [16] to the constrained case (see Theorem 3.8).
  • Under the Bayesian setting, we achieve the same simple regret upper bounds with high probabilities. These bounds are established based on the newly derived bounds (see Theorem 3.13) on the difference between the improvement function and its corresponding EI in the Bayesian setting.
This paper is organized as follows. In Section 2, we describe the basics and preliminaries of BO, including the CEI algorithm. In Section 3, the simple regret upper bounds of CEI are established in both settings. Numerical experiments to validate the theoretical results are given in Section 4. Conclusions are made in Section 5. All proof details are presented in the appendix.
2 Background
CBO consists of two components: the GP surrogates for the black-box objective function $f$ and constraint function $c$ and the constrained acquisition function as the sequential sampling rule guiding for the global optimum.
2.1 Gaussian process models for $f$ and $c$
Without losing generality, let the mean function for the objective GP model prior be 00 and the covariance function (kernel) be $k_{f}({\bm{x}},{\bm{x}}^{\prime}):\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$. At sample point ${\bm{x}}_{t}\in C$, we denote the objective function value as $f({\bm{x}}_{t})$ and the observed constraint function value is $c({\bm{x}}_{t})$. Given $t$ sample points, denote ${\bm{x}}_{1:t}=[{\bm{x}}_{1},\dots,{\bm{x}}_{t}]$ and ${\bm{f}}_{1:t}=[f({\bm{x}}_{1}),\dots,f({\bm{x}}_{t})]$. Moreover, denote the $t\times t$ covariance matrix ${\bm{K}}_{t}^{f}=[k_{f}({\bm{x}}_{1},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{1},{\bm{x}}_{t});\dots;k_{f}({\bm{x}}_{t},{\bm{x}}_{1}),\dots,k_{f}({\bm{x}}_{t},{\bm{x}}_{t})]$. The posterior distribution of $f({\bm{x}})|{\bm{x}}_{1:t},{\bm{f}}_{1:t}\sim\mathcal{N}(\mu_{t}^{f}({\bm{x}}),(\sigma_{t}^{f}({\bm{x}}))^{2})$ can then be inferred using Bayes’ rule as follows
$\mu_{t}^{f}({\bm{x}})\ =\ {\bm{k}}_{t}^{f}({\bm{x}})({\bm{K}}_{t}^{f})^{-1}{\bm{f}}_{1:t},$ (2)
$(\sigma_{t}^{f})^{2}({\bm{x}})\ =\ k^{f}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{f}({\bm{x}}))^{T}({\bm{K}}_{t}^{f})^{-1}{\bm{k}}_{t}^{f}({\bm{x}})\ ,$
where ${\bm{k}}_{t}^{f}({\bm{x}})=[k_{f}({\bm{x}}_{1},{\bm{x}}),\dots,k_{f}({\bm{x}}_{t},{\bm{x}})]^{T}$. Similarly, denote the kernel for $c$ as $k_{c}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ and the covariance matrix ${\bm{K}}_{t}^{c}=[k_{c}({\bm{x}}_{1},{\bm{x}}_{1}),\dots,k_{c}({\bm{x}}_{1},{\bm{x}}_{t});\dots;k_{c}({\bm{x}}_{t},{\bm{x}}_{1}),\dots,k_{c}({\bm{x}}_{t},{\bm{x}}_{t})]$. The posterior distribution for $c$ is
$\mu_{t}^{c}({\bm{x}})\ =\ {\bm{k}}_{t}^{c}({\bm{x}})({\bm{K}}_{t}^{c})^{-1}{\bm{c}}_{1:t},$
$(\sigma_{t}^{c})^{2}({\bm{x}})\ =\ k^{c}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{c}({\bm{x}}))^{T}({\bm{K}}_{t}^{c})^{-1}{\bm{k}}_{t}^{c}({\bm{x}})\ ,$
where ${\bm{k}}_{t}^{c}({\bm{x}})=[k_{c}({\bm{x}}_{1},{\bm{x}}),\dots,k_{c}({\bm{x}}_{t},{\bm{x}})]^{T}$, and $\mu^{c}_{t}({\bm{x}})$ and $(\sigma^{c}_{t})^{2}({\bm{x}})$ are the posterior mean and variance for $c$, respectively. Here we use the superscripts of f and c to distinguish between the GPs for $f$ and $c$.
Choices of the kernels $k_{f}$ and $k_{c}$ include the SE function, Matérn functions, etc. SE and Matérn kernels are among the most popular kernels for GP and BO. Their definitions are as follows.
$k_{SE}({\bm{x}},{\bm{x}}^{\prime})=\exp\left(-\frac{r^{2}}{2l^{2}}\right),\ k_{Mat\acute{e}rn}({\bm{x}},{\bm{x}}^{\prime})=\frac{1}{\Gamma(\nu)2^{\nu-1}}\left(\frac{\sqrt{2\nu}r}{l}\right)^{\nu}B_{\nu}\left(\frac{\sqrt{2\nu}r}{l}\right),$
where $l>0$ is the length hyper-parameters, $r=\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{2}$, $\nu>0$ is the smoothness parameter of the Matérn kernel, and $B_{\nu}$ is the modified Bessel function of the second kind.
2.2 Constrained Expected Improvement
Acquisition functions are critical to the performances of BO algorithms. In the unconstrained setting, one of the most widely adopted acquisition functions is EI [2]. Given $t$ samples, the improvement function of $f$ used in EI is defined as
$I^{f}_{t}({\bm{x}})=\max\{f^{+}_{t}-f({\bm{x}}),0\},$ (3)
where $f^{+}_{t}=\underset{\begin{subarray}{c}i=1,\dots,t\end{subarray}}{\text{min}}\ f({\bm{x}}_{i})$. The expectation of (3) conditioned on existing samples is EI, which has a closed form [19]:
$EI^{f}_{t}({\bm{x}})=(f^{+}_{t}-\mu_{t}^{f}({\bm{x}}))\Phi(z^{f}_{t}({\bm{x}}))+\sigma_{t}^{f}({\bm{x}})\phi(z^{f}_{t}({\bm{x}})),$ (4)
where $z^{f}_{t}({\bm{x}})=\frac{f^{+}_{t}-\mu^{f}_{t}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}$. The functions $\phi$ and $\Phi$ are the probability density function (PDF) and the cumulative distribution function (CDF) of the standard normal distribution, respectively. The $t+1$th sample using EI is chosen by
${\bm{x}}_{t+1}=\underset{\begin{subarray}{c}{\bm{x}}\in C\end{subarray}}{\text{argmax}}EI_{t}^{f}({\bm{x}}).$ (5)
Taking into account the constraint, the constrained improvement function in CEI [1] is defined as
$I^{C}_{t}=\Delta^{c}_{t}({\bm{x}})\max\{f^{+}_{t}-f({\bm{x}}),0\},$ (6)
where $\Delta^{c}_{t}\in\{0,1\}$ is the feasibility indicator function where $\Delta^{c}_{t}({\bm{x}})=1$ if $c({\bm{x}})\leq 0$ and $\Delta^{c}_{t}({\bm{x}})=0$ otherwise. The incumbent $f_{t}^{+}$ in CEI is augmented to be the best feasible observation. CEI assumes that $f$ and $c$ are conditionally independent [1]. Taking the conditional expectation of (6), the CEI function is
$EI^{C}_{t}({\bm{x}})=P_{t}({\bm{x}})EI_{t}^{f}({\bm{x}})=\Phi\left(-\frac{\mu_{t}^{c}({\bm{x}})}{\sigma_{t}^{c}({\bm{x}})}\right)EI_{t}^{f}({\bm{x}}),$ (7)
where $P_{t}(\cdot)$ is the probability of feasibility (POF) function for $c({\bm{x}})\leq 0$. CEI chooses the next sample via
${\bm{x}}_{t+1}=\underset{\begin{subarray}{c}{\bm{x}}\in C\end{subarray}}{\text{argmax}}P_{t}({\bm{x}})EI_{t}^{f}({\bm{x}}).$ (8)
The CEI algorithm is given in Algorithm 1.
Algorithm 1 CEI algorithm
1:Choose $k_{f}(\cdot,\cdot)$, $k_{c}(\cdot,\cdot)$, and $T_{0}$ initial samples ${\bm{x}}_{i},i=1,\dots,T_{0}$. Observe ${\bm{f}}_{1:T_{0}}$ and ${\bm{c}}_{1:T_{0}}$.
2:Train the GP surrogate models for $f$ and $c$ respectively conditioned on the initial observations.
3:for $t=T_{0}+1,T_{0}+2,\dots$ do
4:     Find ${\bm{x}}_{t+1}$ based on (8) (CEI).
5:     Observe $f({\bm{x}}_{t+1})$ and $c({\bm{x}}_{t+1})$.
6:     Update the GP model with the addition of ${\bm{x}}_{t+1}$, $f({\bm{x}}_{t+1})$, and $c({\bm{x}}_{t+1})$.
7:     if Evaluation budget exhausted then
8:         Exit
CEI can be extended to multiple constraints assuming conditional independence among the constraints [1]. Our derived convergence rates can also be readily extended to multiple constraints, as we explain in Remark 3.15.
3 Convergence rates of CEI
We present our main results of convergence rates for CEI by establishing the simple regret upper bounds. Denote the optimal solution to the constrained optimization problem (1) as ${\bm{x}}^{*}$. In the unconstrained case, the simple regret of EI is defined as $f^{+}_{t}-f({\bm{x}}^{*})$ [16]. In the constrained case, we require the current best observation to be feasible when compared to the optimal solution $f({\bm{x}}^{*})$, since one could have an infeasible sample point with smaller objecitve than $f({\bm{x}}^{*})$. Given that $f^{+}_{t}$ is already defined as the best feasible observation till iteration $t$ in CEI, we continue to use
$r_{t}=f^{+}_{t}-f({\bm{x}}^{*}),$ (9)
as the simple regret for CEI.
Remark 3.1 (Infeasible initial samples).
It is well-known that one potential issue with CEI is the requirement of the initial feasible sample [1]. That is, $f_{t}^{+}$ exists from the initial samples so that the CEI calculation can proceed. Recognizing this issue, solutions have been proposed including effective heuristics when $f_{t}^{+}$ does not exist [20] and using the constraint violation function to first find a feasible sample [21]. In our analysis, we make the same underlying assumption as CEI that $f_{t}^{+}$ exists.
We first establish the convergence rate under the frequentist assumptions in Section 3.1, including an improved version of the rate under frequentist assumptions in Section 3.1.1. Then, we establish the convergence rate under Bayesian assumptions in Section 3.2.
3.1 Simple regret upper bound under frequentist assumptions
In this section, we present the simple regret upper bound for CEI under the frequentist setting. Moreover, by adopting the information theory-based bounds and techniques in the noise-free cumulative regret bound of upper confidence bound (UCB) [22], we derive an improved upper bound compared to [16]. The definition of RKHS is given below.
Definition 3.2.
Let $k$ be a positive definite kernel $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ with respect to a finite Borel measure supported on $\mathcal{X}$. A Hilbert space $H_{k}$ of functions on $\mathcal{X}$ with an inner product $\langle\cdot,\cdot\rangle_{H_{k}}$ is called a RKHS with kernel $k$ if $k(\cdot,{\bm{x}})\in H_{k}$ for all ${\bm{x}}\in\mathcal{X}$, and $\langle f,k(\cdot,{\bm{x}})\rangle_{H_{k}}=f({\bm{x}})$ for all ${\bm{x}}\in\mathcal{X},f\in H_{k}$. The induced RKHS norm $\left\lVert{f}\right\rVert_{H_{k}}=\sqrt{\langle f,f\rangle_{H_{k}}}$ measures the smoothness of $f$ with respect to $k$.
In this section, we assume the following assumptions on the functions $f$ and $c$.
Assumption 3.3.
The functions $f$ and $c$ lie in the RKHS, denoted as $\mathcal{H}_{k}^{f}(C)$ and $\mathcal{H}_{k}^{c}(C)$ associated with their respective bounded kernel $k_{f}$ and $k_{c}$, with the norm $\left\lVert{\cdot}\right\rVert_{H_{k}^{f}}$ and $\left\lVert{\cdot}\right\rVert_{H_{k}^{c}}$. The kernels satisfy $k_{f}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, $k_{c}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, $k_{f}({\bm{x}},{\bm{x}})=1$, and $k_{c}({\bm{x}},{\bm{x}})=1$, for $\forall{\bm{x}},{\bm{x}}^{\prime}\in C$. The RKHS norms are bounded above by constants $B_{f}$ and $B_{c}$, respectively, i.e., $\left\lVert{f}\right\rVert_{H_{k}^{f}}\leq B_{f}$, $\left\lVert{c}\right\rVert_{H_{k}^{c}}\leq B_{c}$. Moreover, the bound constraints set $C$ is compact.
Technical Challenges under Frequentist Assumptions.
The main challenge in establishing a simple regret upper bound for CEI is how to incorporate the constraint $c({\bm{x}})\leq 0$ and the probability of feasibility function $P_{t}({\bm{x}})$ into the analysis. Existing regret bounds analysis on CBO methods often focus on UCB-type methods [23, 24], for which the acquisition functions do not have the multiplicative structure between the objective and the constraint, as we pointed out in Section 2.2.
Under Assumption 3.3, both $f$ and $c$ are bounded on $C$ by their RKHS norm bound, as stated in Lemma B.1. The simple regret upper bound is given in the following theorem.
Theorem 3.4.
Under Assumption 3.3, the CEI algorithm leads to the simple regret upper bound of
$r_{t}\leq c_{\tau B}\frac{1}{\Phi(-B_{c})}\left[B_{f}\frac{4}{t-2}+(0.4+B_{f})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (10)
for some $t_{k}\in[\frac{t}{2}-1,t]$, and $c_{\tau B}=\frac{\tau(B_{f})}{\tau(-B_{f})}$.
Sketch of Proof for Theorem 3.4.
We start by noticing that the sum of the difference between consecutive best feasible observations is bounded, i.e., $\sum_{t=1}^{T}f^{+}_{t-1}-f^{+}_{t}\leq 2B_{f}$. Then, we adopt a technique in [16] to find $t_{k}$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, where $k\leq t_{k}\leq 2k$ and $2k\leq t\leq 2(k+1)$. Next, using the monotonicity of $f^{+}_{t}$, $r_{t}$ is bounded by $r_{t_{k}}$. Using the inequality between $I_{t}^{f}$ and $EI_{t}^{f}$ in Lemma B.4, we can bound $r_{t_{k}}$ by the EI on objective: $EI_{t_{k}}^{f}({\bm{x}}^{*})$. Then, we transform $EI_{t_{k}}^{f}({\bm{x}}^{*})$ into $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ by inserting the term $P_{t_{k}}({\bm{x}}^{*})$, taking advantage of the multiplicative structure of CEI. The upper bound of $r_{t}$ then consists of the term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. From the bound on $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ (Lemma B.2) and the fact that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$, we can bound $EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. The constraint term $\frac{1}{P_{t_{k}}({\bm{x}}^{*})}$ remains to be bounded. We use the bound on $|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|$ in Lemma B.2 at ${\bm{x}}^{*}$ and the fact that ${\bm{x}}^{*}$ is a feasible solution to obtain a lower bound for $P_{t_{k}}({\bm{x}}^{*})$. This concludes the proof.
Remark 3.5 (Constraint in the simple regret upper bound).
The terms derived from the constraint function in (10) is $\frac{1}{\Phi(-B_{c})}$, which emerges from the probability of feasibility function and the posterior mean $\mu_{t}^{c}({\bm{x}})$ and standard deviation $\sigma_{t}^{c}({\bm{x}})$ of the GP model of $c({\bm{x}})$. Thanks to the multiplicative structure between the objective and constraint in $I_{t}^{C}$ (6) and $EI_{t}^{c}$ (7), the simple regret upper bound maintains a similar form.
It is clear from (10) that the convergence of $r_{t}$ relies on the posterior standard deviation $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Since $t_{k}$ increases with $t$, as $\sigma_{t}^{f}({\bm{x}}_{t+1})\to 0$, so does $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In the noise-free setting, the posterior variance can be bounded via the maximum distance between sample points and a given point. To obtain the rate of simple regret bound, we use Assumptions (1)-(4) in [16] and focus on squared exponential (SE) and Matérn kernels. Recall that the smoothness parameter of the Matérn kernel is $\nu>0$. Both the SE and Matérn kernels satisfy Assumptions (1)-(4) in [16], with SE kernel obtained as $\nu\to\infty$. Further, define
$\eta=\begin{cases}\alpha,\ \nu\leq 1\\ 0,\ \nu>1,\end{cases}$ (11)
where $\alpha=\frac{1}{2}$ if $\nu\in\mathbb{N}$, and $\alpha=0$ otherwise. Then, for SE and Matérn kernels, $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ can be bounded with the following lemma.
Lemma 3.6 ([16]).
For the SE kernel, there exists constant $C^{\prime}>0$ so that given $\forall t\in\mathbb{N}$,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{1}{d}}$ (12)
holds for at most $k$ times, for $\forall k\in\mathbb{N}$, $k\leq t$, and $i=1,\dots,t-1$. For Matérn kernels,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$ (13)
holds at most $k$ times.
In the constrained setting, we are able to obtain the same rates (14) as those in the unconstrained case [16] using Lemma 3.6.
Theorem 3.7.
Under Assumption 3.3, the CEI algorithm leads to the convergence rates of
$\mathcal{O}\left(t^{-\frac{1}{d}}\right)\ \text{and }\ \mathcal{O}\left(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(t)\right),$ (14)
for SE and Matérn kernels, respectively, where $\eta$ is from (11).
Theorem 3.7 shows that the CEI algorithm is guaranteed to find the best feasible point asymptotically with the rates elaborated in (14). Also, we point out that the choice of kernels and their parameters affect the convergence rates. Since the SE kernel can be viewed as a Matérn kernel with $\nu\to\infty$, its convergence rate is better than Matérn kernels with $\nu\leq 1$. However, due to the limitations of the kernel analysis in [16] (see Remark 3.9), for $\nu\geq 1$, SE and Matérn kernels have similar convergence rates in Theorem 3.7. As we present in the following section, improved rates for both kernels can be obtained.
3.1.1 Improved simple regret upper bound under frequentist assumptions
Next, we apply maximum information gain and the corresponding information theory to obtain improved simple regret upper bounds.
Theorem 3.8.
Under Assumption 3.3, the CEI algorithm leads to the improved convergence rates of
$\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right),$ (15)
for SE and Matérn kernels, respectively.
Sketch of Proof for Theorem 3.8.
The proof follows similar steps to that of Theorem 3.4 but further bounds $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ using $\gamma_{t}^{f}$. To do so, we first recognize that the bound using $\gamma_{t}^{f}$ (Lemma A.3) is established in the noisy case where the posterior standard deviation has a different form as in (19). Using Lemma A.4, we can establish that the noise-free posterior standard deviation also satisfies $\sum_{i=0}^{t-1}\sigma_{i}^{f}({\bm{x}}_{i+1})\leq\sqrt{C_{\gamma}t\gamma_{t}}$. Then, from Lemma A.5, we can find a small enough $\sigma_{i}({\bm{x}}_{i+1})$. Specifically, choose $k=[t/3]$, where $[x]$ denotes the largest integer smaller than $x$. Thus, we have $3k\leq t\leq 3(k+1)$. Then, there exists $k\leq t_{k}\leq 3k$ such that $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}\leq\frac{2B_{f}}{k}$ and $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma_{t}}}{k}$. The rest of the proof follows from that of Theorem 3.8.
Remark 3.9 (Improved rate of convergence).
As mentioned above, the rates in Theorem 3.7 are the same as those to the known convergence rates for EI in [16] with no constraints. Meanwhile, the rates in Theorem 3.8 is an improvement over those of [16] and to the best of our knowledge, are the current best convergence rates for CEI (and EI). To achieve this, we applied techniques from recent regret bound analysis on noise-free UCB [22] that allows us to use maximum information gain to bound the sum of $\sigma_{t}^{f}({\bm{x}}_{t+1})$. Then, we use our techniques in the proof of Theorem 3.4 to bound an individual $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$. In [16], the $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ is bounded by the Taylor expansion of the kernel functions. Therefore, the rates of decrease are limited to quadratic terms for both SE and Matérn kernels, since their Taylor expansions around 00 for $\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{2}$ are quadratic at best. On the other hand, maximum information gain has been a focus of research for years, leading to tight bounds on $\gamma_{t}^{f}$ that take advantages of the spectral properties of the kernels [25]. Hence, using $\gamma_{t}^{f}$ to bound $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ produces a faster rate. As the open question raised in [26] gets answered, further improvement of the convergence rates is possible, e.g., using techniques from [27].
3.2 Simple regret upper bound under Bayesian assumptions
In this section, we present the simple regret upper bound for CEI under the Bayesian assumptions. We again use the maximum information gain to derive the simple regret upper bound.
Assumption 3.10.
The objective function $f$ is sampled from $GP(0,k_{f}({\bm{x}},{\bm{x}}^{\prime}))$. The constraint function $c$ is sampled from $GP(0,k_{c}({\bm{x}},{\bm{x}}^{\prime}))$. The kernels satisfy $k_{f}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, $k_{f}({\bm{x}},{\bm{x}})=1$, $k_{c}({\bm{x}},{\bm{x}}^{\prime})\leq 1$, and $k_{c}({\bm{x}},{\bm{x}})=1$. Moreover, the bound constraint set $C$ is compact.
Denote the probability operator as $\mathbb{P}$. The objective function $f$ is assumed to be Lipschitz continuous under the Bayesian setting, similar to [28].
Assumption 3.11.
There exists constant $L>0$ such that the following holds
$|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert,$ (16)
for $\forall{\bm{x}},{\bm{x}}^{\prime}\in C$.
In the remaining of this section we will work under Assumption 3.10 and 3.11.
Technical Challenges under Bayesian Assumptions.
In addition to the challenges in the frequentist setting, the bounds associated with the Bayesian setting presents further technical challenges, as they are not available in current literature, to the best of our knowledge. Starting from the bounds on $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$, we derive the bounds on $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|$, an important step towards the bound on $r_{t}$ Noticeably, under the Bayesian setting, the bounds are satisfied with a given probability $1-\delta$, where $\delta\in(0,1)$. Hence, the simple regret upper bounds are presented with a probability as well.
The simple regret upper bound is given in the following theorem.
Theorem 3.12.
Let $\beta=2\log(6c_{\alpha}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 3.10 and 3.11, the CEI algorithm leads to the simple regret upper bound
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{4M_{f}}{t-2}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (17)
for some $t_{k}\in[\frac{t}{2}-1,t]$, $c_{\tau}(\beta)=\frac{\tau(\beta^{1/2})}{\tau(-\beta^{1/2})}$, with probability $\geq 1-\delta$.
The convergence rate is given in the next theorem.
Theorem 3.13.
Let $\beta=2\log(6c_{\alpha}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 3.10 and 3.11, the CEI algorithm leads to the convergence rates of
$\mathcal{O}\left(t^{-\frac{1}{2}}\log^{\frac{d+1}{2}}(t)\right)\ \text{and }\ \mathcal{O}\left(t^{\frac{-\nu}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)\right),$ (18)
for SE and Matérn kernels, respectively, with probability $\geq 1-\delta$.
Sketch of Proof for Theorem 3.12.
Recall that $f$ and $c$ are assumed conditionally independent in CEI. We start from the bound on the difference between function value and posterior mean for $f$: $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq\beta^{1/2}\sigma_{t}^{f}({\bm{x}})$, with probability $\geq 1-\delta$, where $\beta=2\log(1/\delta)$, as in Lemma C.2. Similarly, $|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|\leq\beta^{1/2}\sigma_{t}^{c}({\bm{x}})$ with probability $\geq 1-\delta$. These are well-known results [28]. Then, we derive the subsequent bounds $|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})$, where $\beta=\max\{1.44,2\log(c_{\alpha}/\delta)\}$ and $c_{\alpha}=\frac{1+2\pi}{2\pi}$ with probability $\geq 1-\delta$ (Lemma C.4). Then, we prove the relationship in Lemma C.5 that $I_{t}^{f}({\bm{x}})\leq\frac{\tau(\sqrt{\beta})}{\tau(-\sqrt{\beta})}EI_{t}^{f}({\bm{x}})$ with probability $\geq 1-\delta$. We can now follow the general analysis framework in Section 3.1 and Theorem 3.4 to obtain the simple regret upper bound under Bayesian assumptions.
Remark 3.14 (Comparison to the frequentist setting).
Comparing Theorem 3.8 to Theorem 3.13, the convergence rates in the frequentist and Bayesian settings are the same. This is because simple regret focuses on the best feasible solution $f_{t}^{+}$ and thus the parameters in Theorem 3.4 and 3.12 do not depend on $t$. That is, $c_{\tau B}$, $B_{c}$, $B_{f}$, $M_{f}$, and $\beta$ do not depend on $t$. However, we note that $\beta$ depends on the desired probability $\delta$.
Remark 3.15 (Multiple constraints).
As mentioned in Section 1, our results can be readily applied to CEI with multiple constraints for both frequentist and Bayesian settings. Consider $m$ constraints $c_{i}({\bm{x}})\leq 0,i=1,\dots,m$. Assuming conditional independence of the constraints, the CEI function is $EI^{C}_{t}({\bm{x}})=\Pi_{i=1}^{m}P_{t}^{i}({\bm{x}})EI_{t}^{f}({\bm{x}})=\Pi_{i=1}^{m}\Phi\left(-\frac{\mu_{t}^{c_{i}}({\bm{x}})}{\sigma_{t}^{c_{i}}({\bm{x}})}\right)EI_{t}^{f}({\bm{x}})$, where $P_{t}^{i}$ is the probability of feasibility function of constraint $c_{i}({\bm{x}})\leq 0$, and $\mu_{t}^{c_{i}}({\bm{x}})$ and $\sigma_{t}^{c_{i}}({\bm{x}})$ are the posterior mean and standard deviation for $c_{i}$, respectively. By making the assumption that each constraint function lies in its corresponding RKHS of the kernel $k_{c_{i}}$, we have $|c_{i}({\bm{x}})-\mu_{t}^{c_{i}}({\bm{x}})|\leq B_{c_{i}}\sigma_{t}^{c_{i}}({\bm{x}})$, where $B_{c_{i}}$ is the upper bound of RKHS norm associated with kernel $k_{c_{i}}$ and function $c_{i}$. We can then apply the analysis framework in this paper to obtain an upper bound similar to that of Theorem 3.4, where the term $\frac{1}{\Phi(-B_{c})}$ is replaced with $\Pi_{i=1}^{m}\frac{1}{\Phi(-B_{c_{i}})}$.
4 Numerical examples
Although this paper is primarily theoretical, we conduct numerical experiments to support the theoretical results. Specifically, we apply the CEI algorithm on two synthetic examples that are randomly generated and five benchmark problems that are commonly used in the CBO literature. These numerical experiments are not intended to demonstrate superior performance over the state-of-the-art CBO algorithms. Instead, they serve as empirical evidence for the theoretical analysis presented in this work.
We first consider an example where both the objective and constraint functions come from the RKHS of their respective kernels. The functions are generated in a similar approach to [29]. The domain of the problem is $[0,1]^{2}$ and both $f$ and $c$ use SE kernels with a length scale of $0.2$. We randomly generate $100$ basis points, and sample from a multi-variate Gaussian distribution at each point. The posterior means of these samples define the objective and constraint functions, whose contours are shown in Figure 2. This example corresponds to the frequentist setting.
Example 2 considers functions sampled from GP with a Matérn kernel (length scale $0.2$ and $\nu=2.5$), following an approach similar to [28]. We uniformly choose $30\times 30$ points in $[0,1]^{2}$, and sample randomly from a multi-variate zero-mean Gaussian distribution. The contours of the objective and constraint functions are shown in Figure 3. This example corresponds to the Bayesian setting.
(a) Example 1: functions in RKHS of SE kernel.
(b) Example 2: functions sampled from GP with Matérn kernel and $\nu=2.5$
(c) Example 3: a small feasible region problem
(d) Example 4: a multiple-constraints problem
(e) Example 5: a four-dimensional problem
(f) Example 6: Hartmann6 function
(g) Example 7: Banana function (The plot is shown in log scale due to the large range of values.)
Figure 1: Simple regret of CEI for seven examples.
Next, we evaluate simple regret of CEI on five benchmark examples. All these benchmark examples are chosen from the most commonly used test problems in CBO. Specifically, Example 3 tests performance in a small feasible region, which was previously studied in [1, 30]. Example 4 includes multiple constraints and local minimums, which has been used in [31, 32, 33, 30]. Example 5 is a four-dimensional problem, previously studied in [32, 30]. Example 6 is the six-dimensional Hartmann problem, previously tested in [34]. Example 7 is the so-called banana function, where the global minimum lies in a narrow region. The mathematical formulations of Examples 3–7 are presented in Section D in the appendix. For two-dimensional problems, we also include the contour plots of the objective and constraint functions in Appendix D.
Each test problem is repeated for 100 independent experiments. The hyper-parameters of the kernels are estimated by a standard maximum likelihood method. When CEI fails to identify a feasible sample, we adopt the same heuristic strategy as in [34]. The numerical results are summarized in Figure 1(c) - Figure1(g). The solid line represents the median of the simple regret, and the dotted lines represent the 25th percentile and 75th percentile of the simple regret, respectively.
From the figures, we observe that CEI consistently reduces simple regret, aligning with the asymptotic convergence theories established in this paper. For each of the seven examples, the simple regret converges to 00 quickly. The 25th percentile and 75th percentile results demonstrate the good statistical properties of CEI.
5 Conclusions
In this paper, we studied the simple regret upper bounds of the CEI algorithm, one of the most widely adopted CBO methods. Under both frequentist and Bayesian assumptions on the objective and constraint functions, we establish for the first time the convergence rates for CEI. Our results provide theoretical support and validation for the empirical success of CEI.
Appendix A Background information and preliminary results
A.1 Information gain
To obtain state-of-the-art simple regret upper bound, we utilize the information theory results that are well-established in previous literature [28, 29, 25]. Using $f$ as the example, let $A\subset C$ denote a set of sampling points. Assume that the observations are noisy at these points are noisy with $y_{A}=f({\bm{x}}_{A})+\epsilon_{A}$ at ${\bm{x}}\in A$, where $\epsilon_{A}\sim\mathcal{N}(0,\sigma^{2})$ denotes the independent and identically distributed Gaussian noises. The maximum information gain is defined as follows.
Definition A.1.
Given ${\bm{x}}_{A}$ and ${\bm{y}}_{A}$, the mutual information between $f$ and ${\bm{y}}_{A}$ is $I({\bm{y}}_{A};f_{A})=H({\bm{y}}_{A})-H({\bm{y}}_{A}|f_{A})$, where $H$ denotes the entropy. The maximum information gain $\gamma_{T}^{f}$ after $T$ samples is $\gamma_{T}^{f}=\max_{A\subset C,|A|=T}I({\bm{y}}_{A};{\bm{f}}_{A})$.
The rate of increase for $\gamma^{f}_{t}$ depends on the property of the kernel. For common kernels such as the SE kernel and the Matérn kernel, $\gamma^{f}_{t}$ has been widely studied in literature and the state-of-the-art rates of $\gamma^{f}_{t}$ are summarized in Lemma A.2 [25].
Lemma A.2.
For GP with $t$ samples, the SE kernel has $\gamma_{t}=\mathcal{O}(\log^{d+1}(t))$, and the Matérn kernel with smoothness parameter $\nu>0$ has $\gamma_{t}=\mathcal{O}(t^{\frac{d}{2\nu+d}}(\log^{\frac{2\nu}{2\nu+d}}(t)))$.
The maximum information gain $\gamma_{t}^{c}$ for the constraint function can be defined similarly.
While $\gamma_{t}^{f}$ is defined in the noisy case, we can readily apply it to the noise-free case and bound $\sigma_{t}^{f}({\bm{x}}_{t+1})$ using techniques similar to that in [22]. To do so, we note that given the Gaussian observation noise $\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$, we can write the posterior prediction for $f$ as
$\tilde{\mu}_{t}^{f}({\bm{x}})\ =\ {\bm{k}}_{t}^{f}({\bm{x}})({\bm{K}}_{t}^{f}+\sigma^{2}{\bm{I}})^{-1}{\bm{f}}_{1:t},$ (19)
$(\tilde{\sigma}_{t}^{f})^{2}({\bm{x}})\ =\ k^{f}({\bm{x}},{\bm{x}})-({\bm{k}}_{t}^{f})^{T}({\bm{x}})({\bm{K}}_{t}^{f}+\sigma^{2}{\bm{I}})^{-1}{\bm{k}}_{t}^{f}({\bm{x}})\ ,$
Similarly, we can define the posterior predictions for $c$ as $\tilde{\mu}_{t}^{f}({\bm{x}})$ and $\tilde{\sigma}_{t}^{f}({\bm{x}})$. The sum of posterior variance for GP generated by (2) satisfy the next lemma, based on information theory [28] .
Lemma A.3.
The sum of GP posterior variances given $t$ samples satisfy
$\sum_{t=1}^{T}\tilde{\sigma}_{t-1}^{c}({\bm{x}}_{t})\leq\sqrt{C_{\gamma}T\gamma^{c}_{T}},\ \sum_{t=1}^{T}\tilde{\sigma}_{t-1}^{f}({\bm{x}}_{t})\leq\sqrt{C_{\gamma}T\gamma^{f}_{T}},$ (20)
where $C_{\gamma}=\frac{2}{log(1+\sigma^{-2})}$ and $\gamma_{t}^{f}$ and $\gamma_{t}^{c}$ are the maximum information gains for $f$ and $c$, respectively.
We have the following lemma for $\tilde{\sigma}_{t}^{f}({\bm{x}})$ and $\sigma_{t}^{f}({\bm{x}})$.
Lemma A.4.
The noise-free posterior standard deviation satisfies $\sigma_{t}^{f}({\bm{x}})<\tilde{\sigma}_{t}^{f}({\bm{x}})$ for $\forall\sigma>0$.
Proof.
We first note that all the eigenvalues of ${\bm{K}}_{t}^{f}$ is smaller than those of ${\bm{K}}_{t}^{f}+\sigma^{2}{\bm{I}}$, since ${\bm{K}}_{t}^{f}$ is symmetric and positive definite. Thus,
$({\bm{k}}_{t}^{f})^{T}({\bm{x}})({\bm{K}}_{t}^{f}+\sigma^{2}{\bm{I}})^{-1}{\bm{k}}_{t}^{f}({\bm{x}})<({\bm{k}}_{t}^{f})^{T}({\bm{x}})({\bm{K}}_{t}^{f})^{-1}{\bm{k}}_{t}^{f}({\bm{x}}),$ (21)
for $\forall\sigma>0$ and $\forall{\bm{k}}_{t}^{f}({\bm{x}})$. Therefore, by their definitions (2) and (19), the proof is complete. ∎
The posterior standard deviation under assumptions (1)-(4) in [16] in the frequentist and noise-free setting is given in Lemma 3.6, whose proof is given below.
Proof.
Both the SE and Matérn kernels satisfy assumptions (1)-(4) in [16]. By Lemma 7 in [16], there exists $C^{\prime}>0$ so that
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k),$ (22)
at most $k$ times, for $\forall k\in\mathbb{N}$, $k\leq t$, and $i=1,\dots,t-1$. Therefore, for the SE kernel,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{1}{d}},$ (23)
at most $k$ times. For Matérn kernels, we have (22) at most $k$ times. ∎
Using maximum information gain, a tighter bound on the posterior standard deviation can be obtained in the next lemma.
Lemma A.5.
Given $\forall t\in\mathbb{N}$ and $i=1,2,\dots,t-1$, for SE kernel, there exists constant $C^{\prime}>0$ so that
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}\frac{t^{\frac{1}{2}}\log^{\frac{d+1}{2}}(t)}{k},$ (24)
holds for at most $k$ times, for $\forall k\leq t$. Similarly, for Matérn kernels,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}\frac{t^{\frac{\nu+d}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)}{k},$ (25)
holds at most $k$ times.
Proof.
From Lemma A.4 and Lemma A.3, we can write that
$\sum_{i=1}^{t}\sigma_{i-1}^{f}({\bm{x}}_{i})\leq\sqrt{t\gamma_{t}^{f}},$ (26)
where we use without losing generality $C_{\gamma}\leq 1$. Therefore, for any $k\in\mathbb{N}$ and $k\leq t$,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq\frac{\sqrt{t\gamma_{t}^{f}}}{k},$ (27)
at most $k$ times. Therefore, for SE kernel, by Lemma A.2, there exists $C^{\prime}>0$ such that
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}\frac{t^{\frac{1}{2}}\log^{\frac{d+1}{2}}(t)}{k},$ (28)
at most $k$ times. For Matérn kernels,
$\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}\frac{t^{\frac{\nu+d}{2\nu+d}}\log^{\frac{\nu}{2\nu+d}}(t)}{k},$ (29)
at most $k$ times. ∎
The Lipschitz continuity of function $f$ extends to the maximum function $f^{+}=\max\{f,0\}$ in the following lemma.
Lemma A.6.
If a function $f:\mathbb{R}^{n}\to\mathbb{R}$ is Lipschitz continuous on $O\subset\mathbb{R}^{n}$ with Lipschitz constant $L$ in 1-norm, then its maximum with 00 is also Lipschitz, i.e.
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1},$ (30)
where $f^{+}({\bm{x}})=\max\{f({\bm{x}}),0\}$ and ${\bm{x}},{\bm{x}}^{\prime}\in O$.
Proof.
From the definition of Lipschitz functions, we know
$|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1},$ (31)
If $f({\bm{x}})>0$ and $f({\bm{x}}^{\prime})>0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (32)
If $f({\bm{x}})>0$ and $f({\bm{x}}^{\prime})\leq 0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=f({\bm{x}})\leq f({\bm{x}})-f({\bm{x}}^{\prime})\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (33)
Similarly, if $f({\bm{x}})\leq 0$ and $f({\bm{x}}^{\prime})>0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=f({\bm{x}}^{\prime})\leq f({\bm{x}}^{\prime})-f({\bm{x}})\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (34)
Finally if $f({\bm{x}})\leq 0$ and $f({\bm{x}}^{\prime})\leq 0$, then
$|f^{+}({\bm{x}})-f^{+}({\bm{x}}^{\prime})|=0\leq|f({\bm{x}})-f({\bm{x}}^{\prime})|\leq L\left\lVert{{\bm{x}}-{\bm{x}}^{\prime}}\right\rVert_{1}.$ (35)
Next, we state some basic properties of $\phi$, $\Phi$ and $\tau$ as a lemma.
Lemma A.7.
The PDF and CDF of standard normal distribution satisfy $0<\phi(x)\leq\phi(0)<0.4,\Phi(x)\in(0,1)$, for any $x\in\mathbb{R}$. Given a random variable $\xi$ sampled from the standard normal distribution: $\xi\sim\mathcal{N}(0,1)$, we have $\mathbb{P}\{\xi>c|c>0\}\leq\frac{1}{2}e^{-c^{2}/2}$. Similarly, for $c<0$, $\mathbb{P}\{\xi<c|c<0\}\leq\frac{1}{2}e^{-c^{2}/2}$. The function $\tau(\cdot)$ is monotonically increasing.
The last statement in Lemma A.7 is a well-known result (e.g., see proof of Lemma 5.1 in [28]).
The next lemma proves basic properties for $EI_{t}^{f}$.
Lemma A.8.
For $\forall{\bm{x}}\in C$, $EI_{t}^{f}({\bm{x}})\geq 0$ and $EI_{t}^{f}({\bm{x}})\geq y^{+}_{t}-\mu_{t}^{f}({\bm{x}})$. Moreover,
$z_{t}^{f}({\bm{x}})\leq\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}<\begin{cases}\phi(z_{t}^{f}({\bm{x}})),\ &z_{t}^{f}({\bm{x}})<0,\\ z_{t}^{f}({\bm{x}})+\phi(z_{t}^{f}({\bm{x}})),\ &z_{t}^{f}({\bm{x}})\geq 0.\end{cases}$ (36)
Proof.
From the definition of $I_{t}^{f}$ and $EI_{t}^{f}$, $EI_{t}^{f}({\bm{x}})\geq 0$ and $EI_{t}^{f}({\bm{x}})\geq y^{+}_{t}-\mu_{t}^{f}({\bm{x}})$ follow immediately. By (4),
$\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}=z_{t}^{f}({\bm{x}})\Phi(z_{t}^{f}({\bm{x}}))+\phi(z_{t}^{f}({\bm{x}})).$ (37)
If $z_{t}^{f}({\bm{x}})<0$, or equivalently $f_{t}^{+}-\mu_{t}^{f}({\bm{x}})<0$, (37) leads to $\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}<\phi(z_{t}^{f}({\bm{x}}))$. If $z_{t}^{f}({\bm{x}})\geq 0$, we can write $\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}<z_{t}^{f}({\bm{x}})+\phi(z_{t}^{f}({\bm{x}}))$. The left inequality in (36) is an immediate result of $EI_{t}^{f}({\bm{x}})\geq f^{+}_{t}-\mu_{t}^{f}({\bm{x}})$. ∎
Appendix B Proofs for simple regret upper bound under frequentist assumptions
We state the boundedness result as a Lemma for easy reference.
Lemma B.1.
Under Assumption 3.3, $|f({\bm{x}})|\leq B_{f}$ and $|c({\bm{x}})|\leq B_{c}$ for all ${\bm{x}}\in C$.
Proof.
By Assumption 3.3, we can write
$|f({\bm{x}})|\leq\left\lVert{f}\right\rVert_{H_{k}^{f}}k_{f}({\bm{x}},{\bm{x}})\leq B_{f}.$ (38)
Similarly,
$|c({\bm{x}})|\leq\left\lVert{c}\right\rVert_{H_{k}^{c}}k_{c}({\bm{x}},{\bm{x}})\leq B_{c}.$ (39)
The following Lemma is a well-established result [29].
Lemma B.2.
At any given ${\bm{x}}\in C$ and $t\in\mathbb{N}$, the prediction errors satisfy
$|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq B_{f}\sigma_{t}^{f}({\bm{x}}),|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|\leq B_{c}\sigma_{t}^{c}({\bm{x}}).$ (40)
Next, we present a lemma on the relationship between $I_{t}^{f}$ and $EI_{t}^{f}$, previously seen in [16].
Lemma B.3.
At ${\bm{x}}\in C,t\in\mathbb{N}$,
$I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})\leq B_{f}\sigma_{t}^{f}({\bm{x}}).$ (41)
Lemma B.4.
The improvement function $I^{f}_{t}({\bm{x}})$ and $EI^{f}_{t}({\bm{x}})$ satisfy
$I_{t}^{f}({\bm{x}})\leq\frac{\tau(B_{f})}{\tau(-B_{f})}EI^{f}_{t}({\bm{x}}),$ (42)
for $\forall{\bm{x}}\in C$ and $t\geq 1$.
Proof.
If $f^{+}_{t}-f({\bm{x}})\leq 0$, then $I_{t}^{f}({\bm{x}})=0$. Since $EI_{t}^{f}({\bm{x}})\geq 0$, (42) is trivial. If $f^{+}_{t}-f({\bm{x}})>0$, by Lemma B.2,
$f^{+}_{t}-\mu_{t}^{f}({\bm{x}})$ $=f^{+}_{t}-f({\bm{x}})+f({\bm{x}})-\mu_{t}^{f}({\bm{x}})>f({\bm{x}})-\mu_{t}^{f}({\bm{x}})$ (43)
$>-B_{f}\sigma_{t}^{f}({\bm{x}}).$
From the monotonicity of $\tau$, we have
$\tau\left(\frac{f^{+}_{t}-\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)>\tau(-B_{f}),$ (44)
and therefore,
$EI_{t}^{f}({\bm{x}})=\sigma_{t}^{f}({\bm{x}})\tau\left(\frac{f^{+}_{t}-\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)>\tau(-B_{f})\sigma_{t}^{f}({\bm{x}}).$ (45)
From Lemma B.3,
$I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})\leq B_{f}\sigma_{t}^{f}({\bm{x}}).$ (46)
Applying (46) to (45) leads to
$EI_{t}^{f}({\bm{x}})>\frac{\tau(-B_{f})}{B_{f}+\tau(-B_{f})}I_{t}^{f}({\bm{x}})=\frac{\tau(-B_{f})}{\tau(B_{f})}I_{t}^{f}({\bm{x}}).$ (47)
Proof of Theorem 3.4 is given next.
Proof.
From Lemma B.1,
$\sum_{i=0}^{t-1}f^{+}_{i}-f^{+}_{i+1}=f^{+}_{0}-f_{t}^{+}\leq 2B_{f}.$ (48)
Since $f^{+}_{i}-f^{+}_{i+1}\geq 0$, $f^{+}_{i}-f^{+}_{i+1}\geq\frac{2B_{f}}{k}$ at most $k$ times for any $k\in\mathbb{N}$. Further, $f({\bm{x}}_{t})\geq f_{t}^{+}$ for $\forall t\in\mathbb{N}$. Choose $k=[t/2]$, where $[x]$ is the largest integer smaller than $x$ so that $2k\leq t\leq 2(k+1)$. Then, there exists $k\leq t_{k}\leq 2k$ so that $f_{t_{k}}^{+}-f_{t_{k}+1}^{+}<\frac{2B_{f}}{k}$ and $f_{t_{k}+1}^{+}-f({\bm{x}}_{t_{k}+1})\leq 0$.
From Lemma B.4,
$r_{t}=$ $f^{+}_{t}-f({\bm{x}}^{*})\leq f^{+}_{t_{k}}-f({\bm{x}}^{*})\leq I_{t_{k}}^{f}({\bm{x}}^{*})$ (49)
$\leq$ $\frac{\tau(B_{f})}{\tau(-B_{f})}EI_{t_{k}}^{f}({\bm{x}}^{*})=c_{\tau B}\frac{P_{t_{k}}({\bm{x}}^{*})}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}^{*})$
$\leq$ $c_{\tau B}\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}),$
where $c_{\tau B}=\frac{\tau(B_{f})}{\tau(-B_{f})}$. Using $P_{t}({\bm{x}})\leq 1$, (49) implies
$r_{t}\leq$ $\frac{c_{\tau B}}{P_{t_{k}}({\bm{x}}^{*})}\left[(f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))+\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\right]$ (50)
$\leq$ $\frac{c_{\tau B}}{P_{t_{k}}({\bm{x}}^{*})}\left[(f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))+0.4\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$
where the last inequality uses $\phi(\cdot)<0.4$. From Lemma B.2,
$f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})=$ $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}+f^{+}_{t_{k}+1}-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ (51)
$\leq$ $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}+B_{f}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{2B_{f}}{k}+B_{f}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}).$
Using (51) in (50), we have
$r_{t}\leq$ $\frac{c_{\tau B}}{P_{t_{k}}({\bm{x}}^{*})}\left[\frac{2B_{f}}{k}+(B_{f}+0.4)\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (52)
Next, we consider the function $P_{t_{k}}$ at ${\bm{x}}^{*}$. Using the fact that $c({\bm{x}}^{*})\leq 0$, we have by Lemma B.2,
$\mu_{t_{k}}^{c}({\bm{x}}^{*})\leq B_{c}\sigma_{t_{k}}^{c}({\bm{x}}^{*})+c({\bm{x}}^{*})\leq B_{c}\sigma_{t_{k}}^{c}({\bm{x}}^{*}).$ (53)
Thus,
$\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\geq-B_{c}.$ (54)
From the monotonicity of $\Phi$, we have
$P_{t_{k}}({\bm{x}}^{*})=\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi(-B_{c}).$ (55)
Applying (55) to (50), we have
$r_{t}\leq$ $\frac{c_{\tau B}}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+(B_{f}+0.4)\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (56)
As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$. Further, if $\sigma_{t}({\bm{x}}_{t+1})\to 0$, $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\to 0$. ∎
Proof of Theorem 3.7 is presented next.
Proof.
We consider the convergence rate for $r_{t}$ under additional assumptions for the kernel. From Lemma 3.6, for both SE and Matérn kernels, $\sigma_{i}^{f}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$ at most $k$ times for any $k\in\mathbb{N}$ and $i=1,\dots,t$.
Choose $k=[t/3]$ so that $3k\leq t\leq 3(k+1)$. Following the proof of Theorem 3.4, there exists $k\leq t_{k}\leq 3k$ where $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}<\frac{2B_{f}}{k}$, $f({\bm{x}}_{t_{k}+1})\geq f_{t_{k}+1}^{+}$, and $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})<C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$. Similar to (56), we can obtain
$r_{t}\leq$ $\frac{c_{B_{f}}}{\Phi(-B_{c})}\left[\frac{2M_{f}}{k}+(B_{f}+0.4)C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)\right].$ (57)
The convergence rates follow. ∎
B.1 Proofs for improved simple regret upper bound under frequentist assumptions
Proof of Theorem 3.8 is given below.
Proof.
From the proof of Lemma A.5, we know $\sigma_{i}^{f}({\bm{x}}_{i+1})\geq\frac{\sqrt{t\gamma_{t}^{f}}}{k}$ at most $k$ times for any $k\leq t$ and $i=1,\dots,t$.
Choose $k=[t/3]$ so that $3k\leq t\leq 3(k+1)$. Following the proof of Theorem 3.4, there exists $k\leq t_{k}\leq 3k$ where $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}<\frac{2B_{f}}{k}$, $f({\bm{x}}_{t_{k}+1})\geq f_{t_{k}+1}^{+}$, and $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})<3\frac{\sqrt{t\gamma_{t}^{f}}}{t-3}$. Similar to (56), we can obtain
$r_{t}\leq$ $\frac{c_{B_{f}}}{\Phi(-B_{c})}\left[3\frac{2B_{f}}{t-3}+3(B_{f}+0.4)\frac{\sqrt{t\gamma_{t}^{f}}}{t-3}\right].$ (58)
The convergence rates of the simple regret upper bound follow from Lemma A.5. ∎
Appendix C Proofs for simple upper bound regret under Bayesian assumptions
We state the boundedness of $f$ and $c$ as a Lemma for easy reference.
Lemma C.1.
Under Assumption 3.10 and 3.11, there exists $M_{f}>0$ such that $|f({\bm{x}})|\leq M_{f}$. Similarly, there exists $M_{c}>0$ such that $|c({\bm{x}})|\leq M_{c}$.
We recall a well-known result on the bounds of $|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|$ under Assumption 3.10 [28].
Lemma C.2.
Given $\delta\in(0,1)$, let $\beta=2\log(1/\delta)$. For any given ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
$|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}}),$ (59)
holds with probability $\geq 1-\delta$. Similarly, with $\delta\in(0,1)$ and $\beta=2\log(1/\delta)$.
$|c({\bm{x}})-\mu_{t}^{c}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{c}({\bm{x}}),$ (60)
with probability $\geq 1-\delta$.
Proof.
We prove the inequalities for $f$. Under Assumption 3.10, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}({\bm{x}}),\sigma_{t}^{2}({\bm{x}}))$. By Lemma A.7,
$\mathbb{P}\left\{f({\bm{x}})-\mu_{t}^{f}({\bm{x}})>\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})\right\}=1-\Phi\left(\sqrt{\beta}\right)\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$ (61)
Similarly,
$\mathbb{P}\left\{f({\bm{x}})-\mu_{t}^{f}({\bm{x}})<-\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})\right\}\leq\frac{1}{2}e^{-\frac{\beta}{2}}.$ (62)
Thus,
$\mathbb{P}\left\{|f({\bm{x}})-\mu_{t}^{f}({\bm{x}})|<\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})\right\}\geq 1-e^{-\frac{\beta}{2}}.$ (63)
Let $e^{-\frac{\beta}{2}}=\delta$ and (59) is proven. Similarly, (60) can be proven. ∎
The next lemma address $I_{t}^{f}$ under the Bayesian assumption.
Lemma C.3.
Under Assumption 3.10, the probability distribution of $I_{t}$ satisfies
$\mathbb{P}\{I_{t}^{f}({\bm{x}})\leq a\}=\begin{cases}0,\ &a<0,\\ \Phi\left(\frac{a}{\sigma_{t}^{f}({\bm{x}})}-z_{t}^{f}({\bm{x}})\right),\ &a\geq 0.\end{cases}$ (64)
Proof.
Under Assumption 3.10, at a given $t$, $f({\bm{x}})\sim\mathcal{N}(\mu_{t}^{f}({\bm{x}}),\sigma_{t}^{f}({\bm{x}}))$. Since $I_{t}^{f}({\bm{x}})\geq 0$ for all ${\bm{x}}$, (64) follows immediately if $a<0$. For $a\geq 0$,
$\mathbb{P}\{I_{t}^{f}({\bm{x}})\leq a\}=\mathbb{P}\{f^{+}_{t}-f({\bm{x}})\leq a\}=1-\mathbb{P}\{f({\bm{x}})\leq f^{+}_{t}-a\}.$
Using basic properties of the standard normal CDF,
$1-\mathbb{P}\{f({\bm{x}})\leq f^{+}_{t}-a\}=1-\Phi\left(\frac{f_{t}^{+}-a-\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)=\Phi\left(\frac{a-f_{t}^{+}+\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right).$
Next, we present the relationship between $I_{t}^{f}({\bm{x}})$ and $EI_{t}^{f}({\bm{x}})$. Lemma C.4 was previously first presented and proved in [35].
Lemma C.4.
Given $\delta\in(0,1)$, let $\beta=\max\{1.44,2\log(c_{\alpha}/\delta)\}$, where constant $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Under Assumption 3.10, at given ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
$\mathbb{P}\left\{|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})|\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}})\right\}\geq 1-\delta.$ (65)
Proof.
Given a scalar $w>1$, we consider the probabilities
$\mathbb{P}\left\{I_{t}^{f}({\bm{x}})>\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\right\}\quad\text{and}\quad\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\right\}.$ (66)
Consider the first probability in (66). From Lemma A.8, $EI_{t}({\bm{x}})\geq 0$ for $\forall{\bm{x}}$ and $t$. Therefore, $\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})>0$. From Lemma A.8, Lemma C.3, and the monotonicity of $\Phi$, we have
$\mathbb{P}\left\{I_{t}^{f}({\bm{x}})>\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\right\}=$ $1-\Phi\left(\frac{\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})-f^{+}_{t}+\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)$ (67)
$\leq$ $1-\Phi(w)\leq\frac{1}{2}e^{-\frac{w^{2}}{2}},$
where the last inequality is from Lemma A.7.
For the second probability in (66), we further distinguish between two cases. First, consider $-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})<0$. From Lemma C.3,
$\mathbb{P}\left\{I_{t}^{f}({\bm{x}})<-\sigma_{t}^{f}({\bm{x}})w+EI_{t}({\bm{x}})\right\}=0.$ (68)
Second, consider the premise $-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\geq 0$. By Lemma C.3, we have
$\mathbb{P}\left\{I_{t}^{f}({\bm{x}})<-\sigma_{t}^{f}({\bm{x}})w+EI_{t}^{f}({\bm{x}})\right\}=\Phi\left(-w+\frac{EI_{t}^{f}({\bm{x}})-f^{+}_{t}+\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right).$ (69)
To proceed, we show that $f^{+}_{t}-\mu_{t}^{f}({\bm{x}})\geq 0$. Suppose on the contrary, $f^{+}_{t}-\mu_{t}^{f}({\bm{x}})<0$ and thus $z_{t}^{f}({\bm{x}})<0$. From Lemma A.8,
$\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}<\phi(z_{t}^{f}({\bm{x}}))\leq\phi(0)<1\leq w,$ (70)
which contradicts the premise of this case. Thus, we have $f^{+}_{t}-\mu_{t}^{f}({\bm{x}})\geq 0$ (and $z_{t}^{f}({\bm{x}})\geq 0$). From the definition (4), since $\Phi\in(0,1)$,
$\frac{EI_{t}^{f}({\bm{x}})-f^{+}_{t}+\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}=$ $\left[z_{t}^{f}({\bm{x}})\left(\Phi(z_{t}^{f}({\bm{x}}))-1\right)+\phi(z_{t}^{f}({\bm{x}}))\right]<\phi(z^{f}_{t}({\bm{x}})).$ (71)
In addition, by the premise of this case and Lemma A.8,
$w\leq\frac{EI_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\leq z_{t}^{f}({\bm{x}})+\phi(z_{t}^{f}({\bm{x}})).$ (72)
Given that $w>1$ and $\phi(0)\geq\phi(z_{t}^{f}({\bm{x}}))$, we have
$z_{t}^{f}({\bm{x}})+\phi(0)>z_{t}^{f}({\bm{x}})+\phi(z_{t}^{f}({\bm{x}}))>w,\ z_{t}^{f}({\bm{x}})>w-\phi(0)>0.$ (73)
As $z_{t}({\bm{x}})\geq 0$ increases, $\phi(z_{t}({\bm{x}}))>0$ decreases. Thus, we have
$\frac{z_{t}^{f}({\bm{x}})}{\phi(z_{t}^{f}({\bm{x}}))}>\frac{w-\phi(0)}{\phi(w-\phi(0))},\ \phi(z_{t}^{f}({\bm{x}}))<\frac{\phi(w-\phi(0))}{w-\phi(0)}z_{t}^{f}({\bm{x}}).$ (74)
Denote $c_{1}(w)=\frac{w-\phi(0)}{w-\phi(0)+\phi(w-\phi(0))}$. Applying (74) to (72), we obtain
$c_{1}(w)w<z_{t}^{f}({\bm{x}}),\ \ \phi(z_{t}^{f}({\bm{x}}))<\phi(c_{1}(w)w).$ (75)
Applying (75) and (71) to (69), we obtain
$\mathbb{P}\left\{I_{t}^{f}({\bm{x}})<-w\sigma_{t}^{f}({\bm{x}})+EI_{t}^{f}({\bm{x}})\right\}<$ $\Phi\left(-w+\phi(z_{t}^{f}({\bm{x}}))\right)$ (76)
$<$ $\Phi\left(-w+\phi(c_{1}(w)w)\right).$
Notice that $\phi(c_{1}(w)w)<\phi(c_{1}(w))<\phi(c_{1}(w))w$ due to $w>1$. By the definition of $\Phi$ and mean value theorem,
[…]
$\frac{1}{\sqrt{2\pi}}e^{-\frac{1}{2}(w-\phi(c_{1}(w)w))^{2}}\phi(c_{1}(w)w)\leq\Phi(-w)+\frac{1}{2\pi}e^{-\frac{1}{2}((1-\phi(c_{1}(w)))w)^{2}}e^{-\frac{1}{2}(c_{1}(w)w)^{2}}$
$\leq\Phi(-w)+\frac{1}{2\pi}e^{-\frac{1}{2}c_{2}(w)w^{2}}\leq\frac{1}{2}e^{-\frac{1}{2}w^{2}}+\frac{1}{2\pi}e^{-\frac{1}{2}c_{2}(w)w^{2}},$
where $c_{2}(w)=[1-\phi(c_{1}(w))]^{2}+[c_{1}(w)]^{2}$. The last inequality in (77) again uses Lemma A.7. Notice that $c_{2}(w)$ increases with $w$ and for $w\geq 1.2$, $c_{2}(w)>1$. Thus, $e^{-\frac{1}{2}w^{2}}>e^{-\frac{1}{2}c_{2}(w)w^{2}}$ for $w\geq 1.2$, which simplifies (77) to
$\Phi\bigl{(}-w+$ $\phi(c_{1}(w)w)\bigr{)}$ $<c_{\pi 1}e^{-\frac{1}{2}w^{2}}.$ (78)
where $c_{\pi 1}=\frac{1+\pi}{2\pi}$. Therefore, by (76) and (78), if $w\geq 1.2$,
$\mathbb{P}\left\{I_{t}({\bm{x}})<-\sigma_{t}({\bm{x}})w+EI_{t}({\bm{x}})\right\}<c_{\pi 1}e^{-\frac{1}{2}w^{2}}.$ (79)
Combining (79) with (67) and (68), we have
$\mathbb{P}\left\{\left|I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})\right|>w\sigma_{t}^{f}({\bm{x}})\right\}<c_{\alpha}e^{-\frac{1}{2}w^{2}},$ (80)
where $c_{\alpha}=\frac{1+2\pi}{2\pi}$ for $w\geq 1.2$. The probability in (80) monotonically decreases with $w$. Let $\delta=c_{\alpha}e^{-\frac{1}{2}w^{2}}$. Then, taking the logarithm of $\delta$ leads to $\log(\frac{1+2\pi}{2\pi\delta})=\frac{1}{2}w^{2}$. Let $\beta=\max\{w^{2},1.2^{2}\}$, and the proof is complete. ∎
The relationship between $I_{t}^{f}({\bm{x}})$ and $EI_{t}^{f}({\bm{x}})$ under the GP prior assumption is given in the following lemma.
Lemma C.5.
Given $\delta\in(0,1)$, let $\beta=2\log(2c_{\alpha}/\delta)$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. At given ${\bm{x}}\in C$ and $t\in\mathbb{N}$,
$\frac{\tau(-\sqrt{\beta})}{\tau(\sqrt{\beta})}I_{t}^{f}({\bm{x}})\leq EI_{t}^{f}({\bm{x}}),$ (81)
holds with probability $\geq 1-\delta$
Proof.
From Lemma C.2, with probability $\geq 1-\delta$, (59) stands. If $f^{+}_{t}-f({\bm{x}})\leq 0$, then $I_{t}({\bm{x}})=0$. Since $EI_{t}({\bm{x}})\geq 0$, (81) is trivial. If $f^{+}_{t}-f({\bm{x}})>0$, by Lemma C.2,
$f^{+}_{t}-\mu_{t}^{f}({\bm{x}})$ $=f^{+}_{t}-f({\bm{x}})+f({\bm{x}})-\mu_{t}^{f}({\bm{x}})>f({\bm{x}})-\mu_{t}^{f}({\bm{x}})$ (82)
$>-\sqrt{\beta}\sigma_{t}^{f}({\bm{x}}),$
with probability greater than $\geq 1-\delta/2$. From the monotonicity of $\tau$, we have
$\tau\left(\frac{f^{+}_{t}-\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)>\tau(-\sqrt{\beta}),$ (83)
and therefore,
$EI_{t}^{f}({\bm{x}})=\sigma_{t}^{f}({\bm{x}})\tau\left(\frac{f^{+}_{t}-\mu_{t}^{f}({\bm{x}})}{\sigma_{t}^{f}({\bm{x}})}\right)>\tau(-\sqrt{\beta})\sigma_{t}^{f}({\bm{x}}),$ (84)
with probability greater than $21-\delta/$ 2. Using $\delta/2$ in Lemma C.4,
$I_{t}^{f}({\bm{x}})-EI_{t}^{f}({\bm{x}})\leq\sqrt{\beta}\sigma_{t}^{f}({\bm{x}}),$ (85)
with probability $\geq 1-\delta/2$. Applying (85) to (84) with union bound leads to
$EI_{t}^{f}({\bm{x}})>\frac{\tau(-\sqrt{\beta})}{\sqrt{\beta}+\tau(-\sqrt{\beta})}I_{t}^{f}({\bm{x}})=\frac{\tau(-\sqrt{\beta})}{\tau(\sqrt{\beta})}I_{t}^{f}({\bm{x}}),$ (86)
with probability greater than $1-\delta$. ∎
We can now prove Theorem 3.12.
Proof.
From Lemma C.1,
$\sum_{i=0}^{t-1}f^{+}_{i}-f^{+}_{i+1}=f^{+}_{0}-f_{t}^{+}\leq 2M_{f}.$ (87)
Given that $f^{+}_{i}-f^{+}_{i+1}\geq 0$, $f^{+}_{i}-f^{+}_{i+1}\geq\frac{2M_{f}}{k}$ at most $k$ times for any $k\in\mathbb{N}$. Choose $k=[t/2]$, where $[x]$ is the largest integer smaller than $x$ so that $2k\leq t\leq 2(k+1)$. Then, there exists $k\leq t_{k}\leq 2k$ so that $f_{t_{k}}^{+}-f_{t_{k}+1}^{+}<\frac{2M_{f}}{k}$ and $f({\bm{x}}_{t_{k}+1})\geq f_{t_{k}+1}^{+}$.
From Lemma C.5 and $\beta=2\log(6c_{\alpha}/\delta)$, with probability $\geq 1-\delta/3$,
$r_{t}=$ $f^{+}_{t}-f({\bm{x}}^{*})\leq f^{+}_{t_{k}}-f({\bm{x}}^{*})\leq I_{t_{k}}^{f}({\bm{x}}^{*})$ (88)
$\leq$ $\frac{\tau(\beta^{1/2})}{\tau(-\beta^{1/2})}EI_{t_{k}}^{f}({\bm{x}}^{*})=c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}^{*})}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}^{*})\leq c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}EI_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$
$=$ $c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\left[(f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))+\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\right]$
$\leq$ $c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\left[(f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))\Phi(z_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}))+0.4\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$
where the last inequality uses $\phi(\cdot)<0.4$. From Lemma C.2, we have with probability $\geq 1-\delta/3$,
$f^{+}_{t_{k}}-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})=$ $f^{+}_{t_{k}}-f({\bm{x}}_{t_{k}+1})+f({\bm{x}}_{t_{k}+1})-\mu_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})$ (89)
$\leq$ $f^{+}_{t_{k}}-f^{+}_{t_{k}+1}+\beta^{1/2}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{2M_{f}}{k}+\beta^{1/2}\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1}).$
Applying (89) to (88), we have with probability $\geq 1-2\delta/3$,
$r_{t}\leq$ $c_{\tau}(\beta)\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right].$ (90)
Next, we consider the probability function $P_{t_{k}}$ at ${\bm{x}}^{*}$ and ${\bm{x}}_{t_{k}+1}$. Using the fact that $c({\bm{x}}^{*})\leq 0$, we have by Lemma C.2 at ${\bm{x}}^{*}$ and $t_{k}$,
$\mu_{t_{k}}^{c}({\bm{x}}^{*})\leq\sqrt{\beta}\sigma_{t_{k}}^{c}({\bm{x}}^{*})+c({\bm{x}}^{*})\leq\sqrt{\beta}\sigma_{t_{k}}^{c}({\bm{x}}^{*}),$ (91)
with probability $\geq 1-\delta/3$. Thus, with probability $\geq 1-\delta/3$,
$\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\geq-\sqrt{\beta}.$ (92)
From the monotonicity of $\Phi$, we have with probability $\geq 1-\delta/3$,
$\Phi\left(\frac{-\mu_{t_{k}}^{c}({\bm{x}}^{*})}{\sigma_{t_{k}}^{c}({\bm{x}}^{*})}\right)\geq\Phi\left(-\sqrt{\beta}\right),$ (93)
Using (93), the $P_{t_{k}}$ functions have
$\frac{P_{t_{k}}({\bm{x}}_{t_{k}+1})}{P_{t_{k}}({\bm{x}}^{*},\epsilon)}\leq\frac{1}{\Phi(-\sqrt{\beta})},$ (94)
with probability $\geq 1-\delta/3$. Applying (94) to (90), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\right],$ (95)
with probability $\geq 1-\delta$. As $t\to\infty$, $t_{k}\to\infty$ and $k\to\infty$.
The proof of Theorem 3.13 is next.
Proof.
From Lemma A.5, $\sigma_{i}^{f}({\bm{x}}_{i+1})\geq\frac{\sqrt{\gamma_{t}^{f}t}}{k}$, where $i=1,\dots,t-1$, at most $k$ times for any $k\in\mathbb{N}$ and $k\leq t$. Choose $k=[t/3]$ which leads to $3k\leq t\leq 3(k+1)$. There exists $t_{k}$ so that $\sigma_{t_{k}}^{f}({\bm{x}}_{t_{k}+1})\leq\frac{\sqrt{t\gamma^{f}_{t}}}{k}$. Following the proof for (95), we have
$r_{t}\leq$ $c_{\tau}(\beta)\frac{1}{\Phi(-\sqrt{\beta})}\left[\frac{2M_{f}}{k}+(0.4+\beta^{1/2})\frac{\sqrt{t\gamma^{f}_{t}}}{k}\right],$ (96)
with probability $\geq 1-\delta$. Using Lemma A.2, the proof is complete.
Appendix D Example details
The mathematical formulations of the testing problems are given below. The objective, constraint functions and the optimal $f$ of example 3 is given below.
$f({\bm{x}})=\sin(x_{1})+x_{2},$ (97)
$c({\bm{x}})=\sin(x_{1})\sin(x_{2})+0.95\leq 0,$
$x_{i}\in[0,6],i=1,2,$
$f^{*}=0.25.$
The objective, constraint functions and the optimal $f$ of example 4 is given below.
$f(\mathbf{x})=x_{1}+x_{2}$ (98)
$c_{1}({\bm{x}})=-0.5\sin(2\pi(x_{1}^{2}-2x_{2}))-x_{1}-2x_{2}+1.5\leq 0$
$c_{2}({\bm{x}})=x_{1}^{2}+x^{2}_{2}-1.5\leq 0$
$x_{i}\in[0,1],i=1,2$
$f^{*}=0.6.$
The objective, constraint functions and the optimal $f$ of example 5 is given below.
$f(\mathbf{x})=x_{1}+x_{2}+x_{3}+x_{4}$ (99)
$c_{1}=1.1-\sum_{i=1}^{4}E_{i}\exp\left(\sum_{j=1}^{4}-A_{j,i}(x_{j}-P_{j,i})^{2}\right)$
$x_{i}\in[0,1],i=1,\dots,4$
$E=[1,1.2,3,3.2]^{\top}$
$P=\begin{bmatrix}0.131&0.232&0.234&0.404\\ 0.169&0.413&0.145&0.882\\ 0.556&0.830&0.352&0.873\\ 0.012&0.373&0.288&0.574\end{bmatrix}$
$A=\begin{bmatrix}10&0.05&3&17\\ 3&10&3.5&8\\ 17&17&1.7&0.05\\ 3.5&0.1&10&10\end{bmatrix}$
$f^{*}=0.$
The objective, constraint functions and the optimal $f$ of example 6 is given below.
$f(\mathbf{x})=-\sum_{i=1}^{4}\alpha_{i}\exp\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right)$ (100)
$c({\bm{x}})=\sum_{j=1}^{4}x_{j}-3$
$x_{i}\in[0,1],i=1,\dots,6$
$\alpha=[1.0,1.2,3.0,3.2]^{\top}$
$A=\begin{bmatrix}10&3.0&17&3.5&1.7&8.0\\ 0.05&10&17&0.1&8.0&14\\ 3.0&3.5&1.7&10&17&8.0\\ 17&8.0&0.05&10&0.1&14\end{bmatrix}$
$P=\begin{bmatrix}0.131&0.170&0.557&0.012&0.828&0.587\\ 0.233&0.414&0.831&0.374&0.100&0.999\\ 0.235&0.145&0.352&0.288&0.305&0.665\\ 0.405&0.883&0.873&0.574&0.109&0.038\end{bmatrix}$
$f^{*}=-3.32.$
The objective, constraint functions and the optimal $f$ of example 7 is given below.
$f(\mathbf{x})=100(x_{2}-x_{1}^{2})^{2}+(1-x_{1})^{2}$ (101)
$c_{1}({\bm{x}})=\sqrt{x_{1}^{2}+x_{1}^{2}}-4$
$c_{2}({\bm{x}})=x_{1}^{2}+x^{2}_{2}-1.5,$
$x_{1}\in[-5,10],\ x_{2}\in[0,15]$
$f^{*}=0.$
The contour plots of example 1, 2, 3, 4, and 7 are given in Figure 2345, and 6.
Figure 2: Contour plots for the objective function (left) and constraint function (right) for example 1. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 3: Contour plots for the objective function (left) and constraint function (right) for example 2. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 4: Contour plots for the objective function (left) and constraint function (right) for example 3. The feasible region is marked on the plots, along with the global optimum (* sign).
Figure 5: Contour plots for the objective function (left) and the two constraint functions (middle and right) for example 4. The feasible region is marked with black line on the objective contour, along with the global optimum (* sign).
Figure 6: Contour plots for the objective function (left) and constraint function (right) for example 7. The feasible region is marked on the plots, along with the global optimum (* sign).